Up to 1980, medal of Nobel Price was weighing approximately 200 g of 23 carat (95.8% by mass) of gold, and it was changed to 175 g of 18 carat (75% by mass) gold.  How many moles of gold has been saved since then?


1
Expert's answer
2021-10-26T05:37:46-0400

m(pure gold old) = 200*95.8/100 = 191.6g

m(pure gold new) = 175*75/100 = 131.25g

m(saved gold per year) = (191.6-131.25)*5(medals per year) = 301.75g

m(total saved gold) = m(s. gold per year)*40 = 12070g

n(total saved gold) = m(t.s.g.)/M(Au) = 12070/197 = 61.269 moles